OpenStudy (swaqqout_kid1):

Which value of x makes the following equation true? 2(x - 1) = -4x + 28 5 8 9 6 @ttb123456789 @EmmaMink

OpenStudy (anonymous):

5, just plug 5 for x both sides

OpenStudy (kamibug):

First distribute the 2. Then subtract both sides by 28 and then divide both sides by -6.

OpenStudy (emmamink):

Start by substituting in each value for x and see which on fits!

OpenStudy (anonymous):

2x+4x=28+2 6x=30 x=5

OpenStudy (anonymous):

medal?